OpenStudy (anonymous):

what is the measure of each exterior angle of a regular 24-sided polygon?

OpenStudy (anonymous):

360 divided by 24 gives the degrees beyond 180. I'll show you on triangles, squares, and hexagons for an example: equilateral triangles have 60-degree interior angles, which means 300-degree exterior angles, which are 120 degrees more than 180 degrees. 120*3=360. Squares have 90-degree interior angles, which means 270-degree exterior angles, which are 90 degrees more than 180 degrees. 90*4=360. Hexagons have 120-degree interior angles, which means 240-degree exterior angles, which are 60 degrees more than 180 degrees. 60*6=360. I could go on, but here's a picture which shows it more clearly: http://www.geom.uiuc.edu/~dwiggins/pict09C.GIF

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

I can haz medalz?